The height h in feet of a person on a waterslide is modeled by the function h(t)= -0.025t^2-0.5t+50, where t is the time in seconds. At the bottom of the slide, the person lands in a swimming pool. To the nearest tenth of a second, how long does the ride last?
1
Expert's answer
2019-09-27T09:46:01-0400

Time of ride satisfies the condition "h(t)= -0.025t^2-0.5t+50=0". Since the time of ride is not less than ,

the solution "t_1=\frac{0.5+\sqrt{0.5^2+4\cdot 0.025\cdot 50}}{-0.025*2}<0" cannot be the time of ride.

So the time of ride is "t_2=\frac{0.5-\sqrt{0.5^2+4\cdot 0.025\cdot 50}}{-0.025*2}=\frac{-0.5+\sqrt{5.25}}{0.05}\approx 35.8" seconds